OpenStudy (anonymous):

Solve the equation of (¼)log6 (a – 3) – log6 3 = 0. A. 22 B. 64 C. 84 D. 96

OpenStudy (zehanz):

You need two rules for logarithms here:\[k \cdot \log_{a}b =\log_{a}b^k \]and\[\log_{a}b -\log_{a}c=\log_{a}\frac{ b }{ c } \]

Parth (parthkohli):

\[\log_6 a - \log_6 b = \log_6 \left(\dfrac{a}{b}\right)\]So \(\log_6 \left(\dfrac{[a - 3]^{1/4}}{3}\right) =0\) is a start

OpenStudy (anonymous):

ok

Parth (parthkohli):

gotta notice that \(0\) can be written as \( \log_6 1\)